OpenStudy (anonymous):

Use the Laws of Exponents to simplify 3^50 over 3^20 Explain how you found your answer and write the answer as an exponent. @britneyspears

OpenStudy (anonymous):

The rule for division of exponents is \[\frac{ x^a }{ x^b }=x^{(a-b)}\]. The answer using this rule is x^30

OpenStudy (anonymous):

??

OpenStudy (anonymous):

3^50 ÷ 3^20 = 3^(50 - 20) = 3^30 You might think of it as 3 x 3 x 3 x 3 (fifty times) divided by 3 x 3 x 3 x 3 (twenty times). A whole bunch of the three's (top and bottom) are going to cancel out. Try it on a scrap of paper. :)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

You're welcome!! Xo!